### v1

# Efficient evaluation of Bernstein-Bézier coefficients of B-spline basis functions over one knot span

Abstract
New differential-recurrence relations for B-spline basis functions are given. Using these relations, a recursive method for finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span is proposed. The algorithm works for any knot sequence which guarantees that all B-spline functions are at least $C^{0}$-continuous. It has good numerical behavior and has an asymptotically optimal computational complexity.
keywords:
B-spline basis functions, Bernstein-Bézier form, recurrence relations, B-spline curves, B-spline surfaces, de Boor-Cox algorithm.
1 Introduction
[…]
If the boundary knot $t_{n}$ has multiplicity equal to $m+1$, the method is asymptotically optimal. In that case, the computational complexity of the new algorithm is $O(n_{\varepsilon}m^{2})$, where $n_{\varepsilon}$ denotes the number of not-empty knot spans. It means the number of floating point operations is proportional to the number of coefficients which have to be found. If the multiplicity of the knot $t_{n}$ is lower than $m+1$, the computational complexity is increased by at most $O(m^{3})$.
We have also shown that the algorithm is numerically efficient and can be used, for example, in fast rendering of B-spline curves and surfaces. Note that the new approach solves these problems with lower computational complexity than using the famous de Boor-Cox algorithm. Furthermore, when the Bernstein-Bézier form of a B-spline (basis) function is known, one can evaluate it in linear time with respect to its degree by performing the geometric algorithm proposed recently by the authors in [18].
Certainly, the Bernstein-Bézier form of all B-spline functions over all not-trivial knot spans can also be found by other methods (cf., e.g., [1, 2, 3, 4, 5, 8, 15, 16, 17]). However, they have higher computational complexity as they rely on finding all Bernstein-Bézier coefficients of all B-spline basis functions of degrees lower than $m$. For details and the review of the possible solutions, see [7, pp. 3–4].
Let $m,n\in\mathbb{N}$. Suppose the knots $t_{i}$ $(-m\leq i\leq m+n)$ satisfy
$\underbrace{t_{-m}\leq\ldots\leq t_{-1}\leq t_{0}}_{\text{boundary knots}}\leq\underbrace{t_{1}\leq\ldots\leq t_{n-1}}_{\text{inner knots}}\leq\underbrace{t_{n}\leq t_{n+1}\leq\ldots\leq t_{n+m}}_{\text{boundary knots}}\qquad(t_{0}<t_{n}).$ (1.1)
The B-spline (basis) functions $N_{m,-m},N_{m,-m+1},\ldots,N_{m,n-1}$ of degree $m$ over $[t_{0},t_{n}]$ are given by
$N_{mi}(u):=(t_{i+m+1}-t_{i})[t_{i},t_{i+1},\ldots,t_{i+m+1}](t-u)^{m}_{+},$ (1.2)
where the generalized divided difference (see, e.g., [9, §4.2] or [10, p. 7]) acts on the variable $t$, and
$(x-c)^{m}_{+}:=\left\{\begin{array}[]{ll}(x-c)^{m}&(x\geq c),\\ 0&(x<c)\end{array}\right.$ (1.3)
is the truncated power function.
Note that, for $i=0,1,\ldots,n-1$,
$N_{0i}(u)=\left\{\begin{array}[]{ll}1&(u\in[t_{i},t_{i+1})),\\[4.30554pt] 0&\mbox{otherwise},\end{array}\right.$ (1.4)
as well as
$\mbox{supp$(N_{mi})$}=[t_{i},t_{i+m+1}),$ (1.5)
where $-m\leq i<n$.
For more properties of the B-spline basis and for applications of splines in computer-aided geometric design, approximation theory and numerical analysis, see, e.g., [9, 10, 11, 12, 13, 14].
Let the adjusted Bernstein-Bézier basis form of the B-spline function $N_{mi}$ $(-m\leq i<n)$ over a single non-empty knot span $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$ ($j=i,i+1,\ldots,i+m$) be
$N_{mi}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}).$ (1.6)
Here $B^{n}_{i}$ $(0\leq i\leq n;\;n\in\mathbb{N})$ is the $i$th Bernstein (basis) polynomial of degree $n$ given by the formula
$B^{n}_{i}(t):=\binom{n}{i}t^{i}(1-t)^{n-i}.$ (1.7)
In [7] (cf. also [6, §3]), the following problem has been considered.
Problem 1.1.
Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{mi}$ over all non-empty knot spans $[t_{j},t_{j+1})\subset[t_{0},t_{n}]$, i.e., for $j=0,1,\ldots,n-1$ and $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
The proposed solution of Problem 1.1 uses the new differential-recurrence relation for B-spline functions of the same degree. Namely, in [7, Thm 3.1], it was proven that if the knot sequence is clamped, i.e.,
$t_{-m}=t_{-m+1}=\ldots=t_{0}<t_{1}<\ldots<t_{n}=t_{n+1}=\ldots=t_{n+m},$ (1.8)
then
$N_{mi}(u)+\dfrac{t_{i}-u}{m}N_{mi}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$ (1.9)
where $i=-m-1,-m+1,\ldots,n-1$, and a convention is adopted that
$N_{pq}\equiv 0\quad\mbox{for $q<-p$ or $q\geq n$}.$ (1.10)
Next, it has been observed that Eq. (1.9) implies the recurrence relation connecting the coefficients $b_{m,k}^{(i,j)},b_{m,k+1}^{(i,j)}$ and $b_{m,k}^{(i+1,j)},b_{m,k+1}^{(i+1,j)}$. See [7, Thm 4.3].
Using the mentioned facts, the fast and numerically efficient algorithm for solving Problem 1.1 has been proposed. See [7, Thm 4.4, Alg. 4.1 and §6]. Recall that if the boundary knot $t_{n}$ has multiplicity equal to $m+1$ the method is asymptotically optimal — its computational complexity is $O(n_{\varepsilon}m^{2})$, where $n_{\varepsilon}$ is the number of the non-trivial knot spans. Otherwise, the complexity increases by at most $O(m^{3})$.
2 The problem
At the end of our paper [7], we wrote that ”Further research is required to find the application of the new differential-recurrence relation in finding the Bernstein-Bézier coefficients of B-spline basis functions over a single knot span”. The point is that the proposed approach is too expensive if we want to compute the Bernstein-Bézier form of B-spline functions over one fixed knot span only. In short, the main reason is as follows: to find the Bernstein-Bézier coefficients of B-spline functions over non-empty interval $[t_{j},t_{j+1})$ using the recursive algorithm proposed by us, one has to previously compute the Bernstein-Bézier form of B-spline functions over all non-empty knot spans $[t_{n-1},t_{n}),[t_{n-2},t_{n-1}),\ldots,[t_{j+1},t_{j+2})$.
Taking into account the mentioned drawback, the main goal of this article is to propose an optimal method for solving the following problem.
Problem 2.1.
Suppose the knot span $[t_{j},t_{j+1})$ is not empty with fixed $j$ $(0\leq j<n)$. Find the adjusted Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$ $(0\leq k\leq m)$ (cf. (1.6)) of all non-trivial functions $N_{mi}$ over $[t_{j},t_{j+1})$, i.e., for $i=j-m,j-m+1,\ldots,j$ (cf. (1.5)).
Note that to solve Problem 2.1, we have to compute $(m+1)^{2}$ Bernstein-Bézier basis coefficients $b_{m,k}^{(i,j)}$. In Section 4, using the results from §3, we give the recursive algorithm for computing all these quantities in $O(m^{2})$ time, which is asymptotically optimal from the computational complexity standpoint. Certainly, the algorithm is useful, for example, when rendering fragments of a B-spline curve or surface (see also applications considered in [7, §5]).
Let us stress that one can solve Problem 2.1 using the methods proposed, e.g., in [1, 2, 3, 4, 5, 8, 15, 16, 17]). However they have higher computational complexity, i.e., at least $O(m^{3})$.
3 Some new results for B-spline basis functions
We start with proving two new properties of B-spline basis functions which allow us to show that the differential-recurrence relation (1.9) holds even if the knot sequence is not clamped (cf. (1.8)).
Lemma 3.1.
For any knot sequence (1.1), the B-spline functions $N_{mi}$ satisfy the following identities:
$mN_{mi}(u)+(t_{m+i+1}-u)N^{\prime}_{mi}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}},$ (3.1)
$mN_{mi}(u)+(t_{i}-u)N^{\prime}_{mi}(u)=m(t_{m+i+1}-t_{i})\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}},$ (3.2)
where $-m\leq i\leq n-1$ (cf. (1.10)) and we assume that for any quantity $Q$
$\frac{Q}{t_{k}-t_{\ell}}:=0\quad\mbox{if the knots $t_{k}$ and $t_{\ell}$ are equal},$ (3.3)
as well as that $N_{-1,k}\equiv 0$ for all $k$.
Proof.
It is easy to check that the lemma is true if $m\in\{0,1\}$. Let us fix $m>1$ and $i$ $(-m\leq i<n)$. First we prove (3.2) assuming that $t_{m+i+1}\neq t_{i+1}$. From the definitions (1.2) and (1.3), we have
$mN_{mi}(u)$ $=$ $m(t_{m+i+1}-t_{i})[t_{i},t_{i+1},\ldots,t_{i+m+1}](t-u)_{+}^{m}$
$=$ $m(t_{m+i+1}-t_{i})[t_{i},t_{i+1}\ldots,t_{i+m+1}]\left\{(t-u)\cdot(t-u)_{+}^{m-1}\right\}.$
Using the Leibniz‘ formula for the generalized divided differences (see, e.g., [9, Thm 4.4.12]), we obtain
$mN_{mi}(u)$ $=$ $m(t_{m+i+1}-t_{i})\Big{\{}[t_{i}](t-u)\cdot[t_{i},t_{i+1},\ldots,t_{i+m+1}](t-u)_{+}^{m-1}$
$\hskip 85.35826pt+1\cdot[t_{i+1},t_{i+2},\ldots,t_{i+m+1}](t-u)_{+}^{m-1}+0+\ldots+0\Big{\}}$
$=$ $-(t_{i}-u)N^{\prime}_{mi}(u)+m(t_{m+i+1}-t_{i})\frac{N_{m-1,i+1}(u)}{t_{i+m+1}-t_{i+1}}.$
Here the generalized divided difference acts on the variable $t$.
Now, if $t_{m+i+1}=t_{i+1}$ then $[t_{i+1},\ldots,t_{i+m+1}](t-u)_{+}^{m-1}\equiv 0$ (cf. (1.1)) and Eq. (3.2) is also true (cf. (3.3)).
The relation (3.1) can be derived in a similar way. ∎
Remark 3.2.
Observe that Eqs (3.1) and (3.2) can be seen as fundamental properties of B-spline basis functions. Indeed, by adding (3.1) multiplied by $u-t_{i}$ and (3.2) multiplied by $t_{m+i+1}-u$ and next dividing by $m(t_{m+i+1}-t_{i})$, one can recover the famous the de Boor-Cox recurrence relation (see, e.g., [13, Eqs. (2.5)]), i.e.,
$N_{mi}(u)=(u-t_{i})\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}+(t_{m+i+1}-u)\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\qquad(-m\leq i<n).$ (3.4)
We can also easily obtain the following well-known representation of the derivative of $N_{mi}$ in terms of two B-spline basis functions od degree $m-1$ (see, e.g., [13, Eqs. (2.7)]):
$N^{\prime}_{mi}(u)=m\left(\frac{N_{m-1,i}(u)}{t_{m+i}-t_{i}}-\frac{N_{m-1,i+1}(u)}{t_{m+i+1}-t_{i+1}}\right)\qquad(-m\leq i<n).$
It is enough to subtract (3.2) from (3.1) and next divide both sides by $t_{m+i+1}-t_{i}$.
Note that in both cases, we use the conventions (1.10) and (3.3).
Using Lemma 3.1, one can observe that the assumption of [7, Thm 3.1] (cf. (1.8)) can be relaxed. As a consequence, we obtain the following result.
Theorem 3.3.
For any knot sequence (1.1), where no inner knot has multiplicity greater than $m$, the following differential-recurrence relation for B-spline functions of the same degree $m$ holds true:
$N_{mi}(u)+\dfrac{t_{i}-u}{m}N_{mi}^{\prime}(u)=\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}\left(N_{m,i+1}(u)+\dfrac{t_{m+i+2}-u}{m}N_{m,i+1}^{\prime}(u)\right),$
where $i=-m-1,-m+1,\ldots,n-1$ (cf. (1.10)).
Proof.
It is enough to observe that the left-hand side of Eq. (3.2) divided by $m(t_{m+i+1}-t_{i})$ is equal to the left-hand side of (3.1) for $i:=i+1$ divided by $m(t_{m+i+2}-t_{i+1})$. ∎
Remark 3.4.
In the sequel, we assume that no inner knot $t_{1},t_{2},\ldots,t_{n-1}$ has multiplicity greater than $m$. This natural assumption guarantees the B-spline functions‘ continuity in $(t_{0},t_{n})$.
Using Theorem 3.3 and the approach presented in [7, §4.2], one can construct the following recurrence relation for the adjusted Bernstein-Bézier coefficients $b_{m,k}^{(i,j)}$ (see (1.6)) of the B-spline functions (cf. also [7, Thm 4.3]).
Theorem 3.5.
Suppose that the knot sequence has the form (1.1) and the assumption in Remark 3.4 holds true. Let $10\leq j\leq n-$ 1 and $j-m<i<j$. Assume that the interval $[t_{j},t_{j+1})$ is non-empty.
The adjusted Bernstein-Bézier coefficients $b_{m,k}^{(i,j)}$,
$N_{mi}(u)=\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad(t_{j}\leq u<t_{j+1}),$
satisfy the following recurrence relation:
$b_{m,k}^{(i,j)}=\dfrac{t_{j}-t_{i}}{t_{j+1}-t_{i}}b_{m,k+1}^{(i,j)}+\dfrac{v_{mi}}{t_{j+1}-t_{i}}\left((t_{j+1}-t_{m+i+2})b_{m,k}^{(i+1,j)}+(t_{m+i+2}-t_{j})b_{m,k+1}^{(i+1,j)}\right),$ (3.5)
where $k=m-1,m-2,\ldots,0$, and $v_{mi}:=(t_{m+i+1}-t_{i})/(t_{m+i+2}-t_{i+1})$.
4 The solution
Recall that we want to efficiently compute the coefficients $b_{m,k}^{(i,j)}$ for all $k=0,1,\ldots,m$ and all $i=j-m,j-m+1,\ldots,j$, where $10\leq j\leq n-$ 1 is fixed and $t_{j}\neq t_{j+1}$ (see Problem 2.1) assuming that the knot sequence has the form (1.1) and no inner knot has multiplicity greater than $m$ (cf. Remark 3.4).
If $m=0$ the problem is trivial, because $b^{(i,j)}_{0,0}=\delta_{ij}$ (cf. (1.4)). Here $\delta_{ij}$ denotes the Kronecker delta, i.e.,
[…]
Certainly, in general, if $i<j-m$ or $i>j$, we have
$b^{(i,j)}_{m,k}=0\qquad(0\leq k\leq m)$
(cf. (1.5)).
4.1 The main idea
Assume that $m>0$. Let us fix $10\leq j\leq n-$ 1 $(t_{j}\neq t_{j+1})$ and $j-m<i<j$. Supposing that the coefficients $b_{m,k}^{(i+1,j)}$ for all $0\leq k\leq m$ have just been computed and that the value of $b_{m,m}^{(i,j)}$ is known, one can use the recurrence relation (3.5) to compute the coefficients
$b_{m,m-1}^{(i,j)},b_{m,m-2}^{(i,j)},\ldots,b_{m0}^{(i,j)}$
in $O(m)$ time.
It means that to solve Problem 2.1 using Theorem 3.5 in $O(m^{2})$ time — i.e., to propose the asymptotically optimal algorithm of computing the adjusted Bernstein-Bézier forms of all non-trivial $N_{mi}$ over non-empty knot span $[t_{j},t_{j+1})$ for all $j-m\leq i\leq j$ — first, one needs to:
  1. 1.
    find the adjusted Bernstein-Bézier forms of $N_{m,j-m}$ and $N_{mj}$,
  2. 2.
    evaluate the quantities $b_{m,m}^{(i,j)}$ for all $i=j-m+1,j-m+2,\ldots,j-1$
in at most $O(m^{2})$ time. We will discuss these issues in the next subsection.
4.2 The initial conditions
Using (3.4) and (1.5), it is not difficult to check that for $m>0$ and any knot sequence (1.1), we have
$N_{m,j-m}(u)=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+1}-t_{j+1-k})}B^{m}_{0}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right),$
$N_{mj}(u)=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+k}-t_{j})}B^{m}_{m}\left(\dfrac{u-t_{j}}{t_{j+1}-t_{j}}\right)\qquad\qquad(t_{j}\leq u<t_{j+1}),$
where $10\leq j\leq n-$ 1 and $t_{j}\neq t_{j+1}$.
It means that the Bernstein-Bézier coefficients of $N_{m,j-m}$ and $N_{mj}$ $(m>0)$ in the non-empty knot span $[t_{j},t_{j+1})$ $(0\leq j\leq n-1)$ are as follows:
$b_{m0}^{(j-m,j)}=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+1}-t_{j-k+1})},\qquad b_{m,k}^{(j-m,j)}=0\quad(k=1,2,\ldots,m),$ (4.2)
$b_{m,k}^{(j,j)}=0\quad(k=0,1,\ldots,m-1),\qquad b_{m,m}^{(j,j)}=\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+k}-t_{j})}$ (4.3)
(cf. [7, §4.1]) and all these coefficients can be computed in total $O(m)$ time.
Now, we show that the quantities $b^{(i,j)}_{m,m}$ satisfy a simple recurrence relation with respect to $m$ and $i$.
Theorem 4.1.
For $m>0$ and any knot sequence (1.1), where no inner knot has multiplicity greater than $m$, the following recurrence relation holds true:
$b_{m,m}^{(i,j)}=\dfrac{t_{j+1}-t_{i}}{t_{m+i}-t_{i}}b_{m-1,m-1}^{(i,j)}+\dfrac{t_{m+i+1}-t_{j+1}}{t_{m+i+1}-t_{i+1}}b_{m-1,m-1}^{(i+1,j)},$ (4.4)
where $10\leq j\leq n-$ 1 is fixed, $t_{j}\neq t_{j+1}$, and $i=j-m,j-m+1,\ldots,j-1$.
Proof.
In the proof we use the approach similar to the one presented in [5, 15].
It is well-known that Bernstein polynomials (1.7) satisfy
$tB^{n}_{k}(t)$ $=$ $\dfrac{k+1}{n+1}B^{n+1}_{k+1}(t),$ (4.5)
$B^{n}_{k}(t)$ $=$ $\dfrac{n-k+1}{n+1}B^{n+1}_{k}(t)+\dfrac{k+1}{n+1}B^{n+1}_{k+1}(t),$ (4.6)
where $0\leq k\leq n$.
Let $u\in[t_{j},t_{j+1})$, where $10\leq j\leq n-$ 1 and $t_{j}\neq t_{j+1}$. Set
$t\equiv t^{(j)}(u):=\dfrac{u-t_{j}}{t_{j+1}-t_{j}}$
(cf. (1.6)). Certainly, $t\in[0,1)$, and
$u=(t_{j+1}-t_{j})t+t_{j}.$ (4.7)
Assume $i=j-m,j-m+1,\ldots,j-1$ (cf. (1.5)). Applying (1.6) to (3.4) gives
$\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}(t)=\dfrac{u-t_{i}}{t_{m+i}-t_{i}}\sum_{k=0}^{m-1}b^{(i,j)}_{m-1,k}B^{m-1}_{k}(t)+\dfrac{t_{m+i+1}-u}{t_{m+i+1}-t_{i+1}}\sum_{k=0}^{m-1}b^{(i+1,j)}_{m-1,k}B^{m-1}_{k}(t).$
Using (4.7) gives, after some algebra,
$\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}(t)=\sum_{k=0}^{m-1}\left(\dfrac{b^{(i,j)}_{m-1,k}}{t_{m+i}-t_{i}}-\dfrac{b^{(i+1,j)}_{m-1,k}}{t_{m+i+1}-t_{i+1}}\right)(t_{j+1}-t_{j})tB^{m-1}_{k}(t)\\ +\sum_{k=0}^{m-1}\left(\dfrac{b^{(i,j)}_{m-1,k}(t_{j}-t_{i})}{t_{m+i}-t_{i}}+\dfrac{b^{(i+1,j)}_{m-1,k}(t_{m+i+1}-t_{j})}{t_{m+i+1}-t_{i+1}}\right)B^{m-1}_{k}(t).$
Applying (4.6) and (4.5) gives, then, after some more algebra,
$\sum_{k=0}^{m}b^{(i,j)}_{m,k}B^{m}_{k}(t)=\sum_{k=0}^{m-1}\left(\dfrac{b^{(i,j)}_{m-1,k}(t_{j}-t_{i})}{t_{m+i}-t_{i}}+\dfrac{b^{(i+1,j)}_{m-1,k}(t_{m+i+1}-t_{j})}{t_{m+i+1}-t_{i+1}}\right)\dfrac{m-k}{m}B^{m}_{k}(t)\\ +\sum_{k=1}^{m}\left(\dfrac{b^{(i,j)}_{m-1,k-1}(t_{j+1}-t_{i})}{t_{m+i}-t_{i}}+\dfrac{b^{(i+1,j)}_{m-1,k-1}(t_{m+i+1}-t_{j+1})}{t_{m+i+1}-t_{i+1}}\right)\dfrac{k}{m}B^{m}_{k}(t).$ (4.8)
Matching the coefficients for $B^{m}_{m}(t)$ justifies the recurrence relation (4.4). ∎
Observe that it is possible to evaluate the quantities $b^{(i,j)}_{m,m}$ for $m>1$ and all $i=j-m+1,j-m+2,\ldots,j-1$ in $O(m^{2})$ time using the recurrence relation (4.4), with the initial conditions
$b_{p,p}^{(i,j)}:=0\quad(i<j-p),\qquad b_{0,0}^{(i,j)}:=\delta_{ij},$ (4.9)
$b_{1,1}^{(j,j)}:=1,\qquad b_{p,p}^{(j,j)}:=\dfrac{t_{j+1}-t_{j}}{t_{j+p}-t_{j}}\cdot b_{p-1,p-1}^{(j,j)}\quad(p>0)$ (4.10)
(see (1.5), (4.1), and (4.3)).
Remark 4.2.
From Eq. (4.8), it follows that
$b^{(i,j)}_{m,k}$ $=$ $\dfrac{m-k}{m}\left(\dfrac{b^{(i,j)}_{m-1,k}(t_{j}-t_{i})}{t_{m+i}-t_{i}}+\dfrac{b^{(i+1,j)}_{m-1,k}(t_{m+i+1}-t_{j})}{t_{m+i+1}-t_{i+1}}\right)$
$\hskip 85.35826pt+\dfrac{k}{m}\left(\dfrac{b^{(i,j)}_{m-1,k-1}(t_{j+1}-t_{i})}{t_{m+i}-t_{i}}+\dfrac{b^{(i+1,j)}_{m-1,k-1}(t_{m+i+1}-t_{j+1})}{t_{m+i+1}-t_{i+1}}\right),$
where $0\leq k\leq m$ (cf. the convention (3.3)). Observe that one can use this recurrence scheme to solve Problem 2.1 in $O(m^{3})$ time.
4.3 The algorithm
The results presented in Theorem 3.5, as well as in §4.1 and §4.2 can be combined to prove the following theorem.
Theorem 4.3.
Suppose that $m>0$, the knot sequence has the form (1.1) and no inner knot has multiplicity greater than $m$. Let us fix $j$ $(0\leq j<n)$. Assume that knot span $[t_{j},t_{j+1})$ is non-empty.
The $(m+1)^{2}$ adjusted Bernstein-Bézier coefficients $b_{m,k}^{(i,j)}$ (cf. (1.6)) of the basis B-spline functions $N_{mi}$ (1.2) over the knot span $[t_{j},t_{j+1})$, for $i=j-m,j-m+1,\ldots,j$ and $k=0,1,\ldots,m$, can be computed in the computational complexity $O(m^{2})$ in the following way:
  1. 1.
    Set $b_{0,0}^{(j,j)}=1$. Then, for $p=1,2,\ldots,m$, compute $b_{p,p}^{(j,j)}=\dfrac{t_{j+1}-t_{j}}{t_{j+p}-t_{j}}b_{p-1,p-1}^{(j,j)}$.
  2. 2.
    For $p=1,2,\ldots,m$ and $i=j-1,j-2,\ldots,j-p$ compute $b_{p,p}^{(i,j)}$ using Eqs (4.4), (4.9) and the convention (3.3).
  3. 3.
    For $k=0,1,2,\ldots,m$, compute $b_{m,k}^{(j-m,j)}$ using Eq. (4.2).
  4. 4.
    For $k=m-1,m-2,\ldots,0$ and $i=j-1,j-2,\ldots,j-m+1$, compute the coefficients $b_{m,k}^{(i,j)}$ using Eq. (3.5).
[…]
Figure 4.1: An illustration of using Eq. (4.4) (cf. (4.9)) for $m=5$.
Figure 4.1 illustrates the computations done in the first two steps of Theorem 4.3. The rightmost column is then inserted into the rightmost column of Figure 4.2, which illustrates the remaining steps.
[…]
Figure 4.2: An illustration of the full recurrence scheme using Theorem 3.5 for $m=5$.
Algorithm 4.1 implements the approach proposed in Theorem 4.3. This algorithm returns an array $B\equiv B[0..m,j-m..j]$, where
[…]
(cf. (1.6)).
The complexity of Algorithm 4.1 is $O(m^{2})$ — giving the optimal $O(1)$ time per coefficient.
To examine the numerical behavior of the new algorithm (complexity $O(m^{2})$), we have compared it to the method mentioned in Remark 4.2 (complexity $O(m^{3})$). Numerical experiments show that both methods are stable and give almost the same results. The source code in C which was used to perform the tests is available at https://www.ii.uni.wroc.pl/~pwo/programs/BSpline-BF-one-interval.c.
Algorithm 4.1 Computing the coefficients of the adjusted Bernstein-Bézier form of the B-spline functions over one knot span
1:procedure BSplineBBFOneKnotSpan($n,m,j,[t_{-m},t_{-m+1},\ldots,t_{n+m}]$)
[…]
3:     $Bm[0,j]\leftarrow 1$
4:     for $p=1,m$ do
5:         $Bm[p,j]\leftarrow\dfrac{t_{j+1}-t_{j}}{t_{j+p}-t_{j}}\times Bm[p-1,j]$
6:     end for
7:     for $p=1,m$ do
8:         for $i=j-1,j-p$ do
9:              // With the convention (3.3).
10:              $Bm[p,i]\leftarrow\dfrac{t_{j+1}-t_{i}}{t_{p+i}-t_{i}}\times Bm[p-1,i]+\dfrac{t_{p+i+1}-t_{j+1}}{t_{p+i+1}-t_{i+1}}\times Bm[p-1,i+1]$
11:         end for
12:     end for
[…]
14:     $B[0,j-m]\leftarrow\dfrac{(t_{j+1}-t_{j})^{m-1}}{\prod_{k=2}^{m}(t_{j+1}-t_{j-k+1})}$
15:     $B[m,j-m..j]\leftarrow Bm[m,j-m..j]$
16:     for $k=m-1,0$ do
17:         for $i=j-1,j-m+1$ do
18:              $v_{mi}\leftarrow\dfrac{t_{m+i+1}-t_{i}}{t_{m+i+2}-t_{i+1}}$
19:              $c_{i}\leftarrow(t_{j+1}-t_{m+i+2})B[k,i+1]+(t_{m+i+2}-t_{j})B[k+1,i+1]$
20:              $B[k,i]\leftarrow\dfrac{t_{j}-t_{i}}{t_{j+1}-t_{i}}B[k+1,i]+\dfrac{v_{mi}}{t_{j+1}-t_{i}}c_{i}$
21:         end for
22:     end for
23:     return $B$
24:end procedure
5 Conclusions
Two new differential-recurrence relations for the B-spline functions are given. They can be used to relax the assumptions (1.8) about the knot sequence and give an alternative proof to the differential-recurrence relation (1.9) for the B-spline functions of the same degree which was given in [7], along with an algorithm for computing the coefficients of B-spline functions of degree $m$ in the adjusted Bernstein-Bézier form. The algorithm [7, Thm 4.4, Alg. 4.1 and §6] had a drawback — the requirement to evaluate the coefficients over multiple knot spans. We have eliminated this constraint, thus making the computations independent for each knot span.
The new algorithm given in this manuscript is asymptotically optimal — computing $(m+1)^{2}$ coefficients in $O(m^{2})$ time — and allows the user to evaluate the coefficients only in the necessary knot spans. Moreover, if needed, the computations for multiple knots spans can be done in parallel, as they are fully independent. The method has good numerical behavior.
When the Bernstein-Bézier coefficients of B-spline basis functions are known, it is possible to compute any B-spline function in linear time with respect to its degree by performing the geometric algorithm proposed recently in [18]. As shown in [7], this can accelerate the evaluation when multiple B-spline curves or surfaces use the same knot sequence.